If the charge is $16 for t=0 minutes and c=$19.50 for 35 minutes, find the linear equation that expresses c in terms of t and use it to determine when t=75 minutes
Answer by lwsshak3(11628)

If the charge is $16 for t=0 minutes and c=$19.50 for 35 minutes, find the linear equation that expresses c in terms of t and use it to determine when t=75 minutes
**
c=mt+b
given two points to work with (16,0) and (19.50,35)
m=∆c/∆t=(19.5-16)/35-0)=3.5/35=1/10
c=t/10+b
find b using one of the given points (16,0)
16=0/10+b
b=16
Equation:
c=t/10+16
for t=75 min
c=75/10+16=$23.50